OpenStudy (anonymous):

PLZ help!! medal for correct answer! what is the value of Z so that -9 and 9 are both solutions of x^2+z=103

OpenStudy (anonymous):

@nashg you know?

OpenStudy (anonymous):

this one to

OpenStudy (anonymous):

i got this first question i just need the one on the pic i posted, i aslo posted the wrong answers forthe first one

OpenStudy (anonymous):

forget the first question xD

OpenStudy (campbell_st):

well you need the difference of 2 squares to get the solutions of -9 and 9 so the equation would be \[x^2 - 81 = 0~~~or~~~~~x^2 = 81\] so by subtracting z from both sides you need to solve \[81 = 103 - z\]

OpenStudy (campbell_st):

and just solve for z

OpenStudy (anonymous):

right, i got 22 right?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

is number 4 no solution?

OpenStudy (campbell_st):

that's correct for z.

OpenStudy (anonymous):

ok cool. i assume that because i plugged them in and it dosent make sense

OpenStudy (campbell_st):

well you know the area is 24 and side lengths are 3b and b so then using the formula for area of a triangle \[24 = \frac{1}{2} 3b \times b\] you can solve for b. One suggestion is that measurements can only be positive.... so that may help you get the correct answer

OpenStudy (anonymous):

so do i plug in answers to see it they fit?

OpenStudy (anonymous):

iits 4? cuz 3*4=12 half of 12 is 6 nad 6*4 is 24 right?